a piece of plywood, the surface of which measures eight-ninths square yard, is cut into sections each containing two-ninths square yard. how many sections are there?
Answer by mananth(16949)

the surface of the plywood = 8/9 sq. yards
the area of cut portions =2/9 sq. yards
Number of pieces = 8/9 / 2/9
= 8/9 * 9/2
4 pieces